OpenStudy (cheesecakekitten):

PLEASE HELP ME I CAN'T FAIL 9TH GRADE

OpenStudy (cheesecakekitten):

Find a single expression that represents the surface area of the rectangular prism. Rectangular Prism: http://s30.postimg.org/ggpzc51q9/Even_More_Math_Junk_x1107.png A) 9x2 + 4 B) 12x2 + 20 C) 14x2 + 12 D) 18x2 + 8

OpenStudy (lpruitt4):

9x2+2

OpenStudy (cheesecakekitten):

more plz

OpenStudy (cheesecakekitten):

oh wait

OpenStudy (cheesecakekitten):

that answer was wrong ;-;

OpenStudy (cheesecakekitten):

oh well next question

OpenStudy (lpruitt4):

that is the right one

OpenStudy (cheesecakekitten):

Multiply: (2x2 − 5x)(3x2 + 9x + 3) A) 6x4 + 3x3 − 39x2 − 15x B) 6x4 + 3x3 − 51x2 − 15x C) 6x4 − 33x3 + 51x2 − 15x D) 6x4 − 33x3 + 51x2 + 15x

OpenStudy (lpruitt4):

a or d cuz it make since

OpenStudy (cheesecakekitten):

a

OpenStudy (lpruitt4):

yes

OpenStudy (lpruitt4):

is it right

OpenStudy (cheesecakekitten):

YAP

OpenStudy (cheesecakekitten):

Andy and Sam are saving money to go on their senior trip. The amount of money that Andy will have at the end of each week, w, can be expressed at A(w) = 20w + 300. The amount of money that Sam will have at the end of each week, w, can be expressed at S(w) = 15w + 400. They have decided to combine their savings accounts. Write a function that expresses the total amount, T(w) they have in their savings account at the end of each week. A) T(w) = 5w - 100 B) T(w) = 35w + 700 C) T(w) = -5w + 100 D) T(w) = 300w^2 + 12500w + 120000

OpenStudy (lpruitt4):

at the end i will tell you a big sercet

OpenStudy (cheesecakekitten):

oh okay but can you help plz

OpenStudy (cheesecakekitten):

@lpruitt4

OpenStudy (cheesecakekitten):

if you don't get back here i'm giving your medal to someone else

OpenStudy (cheesecakekitten):

OKay, who wants a medal?